# POLAR: A Benchmark for Multilingual, Multicultural, and Multi-Event Online Polarization

### v2

@@ 3.3 Annotation Process
  • If the text is classified as polarized, tag its manifestations (stereotyping/deindividuation, vilification, dehumanization, extreme language, lack of empathy, invalidation).
@@
Annotator Recruitment and Annotation Process: We used both expert and crowd-sourced annotation strategies. Trained annotators used POTATO and Label Studio, while crowdsourced annotators used Mechanical Turk and Prolific. To evaluate inter-annotator agreement, we report both Cohen’s Kappa and Fleiss’ Kappa. Since different teams were responsible for dataset creation across various languages, we provide a detailed description of the annotation process for each language:
  • Amharic: We used trained annotators to annotated 5,000 samples in five batches using the POTATO annotation tool. Annotators completed three rounds of guideline training and received continuous feedback, with ongoing supervision to ensure quality control. The annotation quality achieved a Fleiss’ Kappa of 0.49, with full agreement in 69.62% of cases and a peak pairwise Cohen’s Kappa of 0.65.
  • Arabic: Three topic-aware annotators labeled 2,118 samples using POTATO. Final polarization distribution was 16% positive, with Cohen’s Kappa at 0.296. Annotations were revised to ensure thematic consistency.
  • German: Annotation was conducted through Prolific. Annotators were screened via surveys and underwent pilot testing. Final annotation used a locally hosted POTATO interface adapted for multilabel inputs. Inter-annotator agreement was moderate, with notable consistency on polarized cases.
  • Hausa: We used trained annotators and Label Studio with iterative feedback and internal consistency checks. Data focused on religion and ideology on social media.
  • Spanish and English: Amazon Mechanical Turk (MTurk) was used with strict quality filters, such as task completion history and approval ratings. Annotators received feedback and were contacted throughout for quality assurance. English annotation reached a Kappa of 0.52 (up from 0.31 after worker filtering); Spanish showed lower agreement with an average Kappa of 0.24.
  • Urdu: Two approaches were used: (1) Manual Annotation: 1,792 samples labeled in batches with average Cohen’s Kappa of 0.55 (batch 1: 0.79; batch 2: 0.56; batch 3: 0.39). Disagreements were resolved by a fifth annotator. (2) Prolific: 1,460 samples were annotated in three batches. Initial agreement was low (Fleiss’ Kappa = 0.05). Additional filters (education, region) raised agreement in later batches to 0.28 and 0.30. The final dataset contains 2,800 high-quality Urdu samples.
Figure 3: Percentage of polarization per language
3.4 Dataset Characteristics
Language Total Polarization Polarization Types Polarization Manifestations
gender/sexual political religious racial/ethnic other vilification extreme_language stereotype invalidation lack_of_empathy dehumanization
Amharic 5000 3753 29 3342 99 1297 189 2398 1527 2729 799 880 657
Arabic 2118 338 102 178 86 171 220 413 341 283 240 175 102
English 3025 1021 18 892 44 98 5 342 179 138 85 45 39
German 3467 1596 241 1981 448 782 653 1371 1001 1420 1258 1124 568
Hausa 5563 625 46 285 149 183 22 74 179 253 13 53 202
Spanish 2002 1057 254 742 571 521 405 798 622 843 205 599 307
Urdu 2800 1860 236 1533 557 423 126 1548 1409 979 670 632 501
Table 2: Number of samples labeled positive for each annotation task across languages. Labels are grouped by task: Polarization (Task 1), Polarization Types (Task 2), and Polarization Manifestations (Task 3).
We now analyze key trends and insights across the dataset based on annotated labels. Table 2 provides a detailed quantitative breakdown of positive labels for each annotation task.
Polarization Prevalence: As shown in Figure 3 and further detailed in Table 2, Amharic (75%) and Urdu (66%) have the highest polarization rates, tied to their sources - Northern Ethiopia’s Tigray conflict and Pakistan’s political and sectarian climate. Spanish (53%), German (46%), and English (34%) fall in the mid-range. Arabic (12%) and Hausa (11%) reflect broader topical focus and possibly platform moderation effects. Hausa shows the lowest polarization rate despite covering sensitive religious and ideological topics, possibly due to moderation or cultural norms in expression. Datasets with acute sociopolitical events (e.g., Amharic, Urdu) show higher polarization. Topic scope, moderation, and platform dynamics significantly influence polarization prevalence.
Types of Polarization: As illustrated in Figure LABEL:fig:types and Table 2, political polarization dominates in Amharic (67%), German (57%), Urdu (55%), and Spanish (37%). Racial/ethnic and religious types also appear prominently in Spanish, Amharic, and German. Gender/sexual identity polarization is generally low but visible in Spanish (13%) and Urdu (8%). Thematic emphasis (e.g., migration, elections) shapes type distribution. Spanish and Urdu show diverse forms of identity-based polarization, while political conflict dominates Amharic, Urdu, and German.
Polarization Manifestations: Figure LABEL:fig:manifestation and Table 2 show the types of manifestations. Stereotyping/deindividuation and vilification are most frequent in Urdu, Amharic, German, and Spanish. Urdu’s vilification rate is highest (55%). Extreme language is strong in Urdu (50%), Spanish (31%), and Amharic (31%). Invalidation and dehumanization peak in German and Urdu. Conflict-heavy datasets (Urdu, Amharic) correlate with higher rates of hostile manifestations. Platform moderation and topic scope likely suppress such features in Arabic and Hausa.
Together, these analyses highlight how linguistic, cultural, and contextual variables - alongside event salience and platform dynamics - shape the structure and tone of online polarization. The POLAR dataset offers a robust benchmark for examining these effects at scale.
@@ 4.1 Experimental Setup
To evaluate POLAR , we conducted baseline experiments on three polarization detection tasks:
  • Task 1: Binary classification - determining whether a text is polarized.
  • Task 2: Multi-label classification - identifying polarization types (e.g., political, religious, ethnic).
  • Task 3: Multi-label classification - detecting polarization manifestations (e.g., stereotyping, vilification, dehumanization).
For data split, we used 70% for training, 10% for validation, and 20% for testing, as summarized in Table˜1. We pursued two main experimental paradigms:
  1. 1.
    Fine-tuning Multilingual Language Models (MLMs): We fine-tuned six multilingual models including InfoXLM (chi2021infosxlm), LaBSE (feng2022language), RemBERT (chung2020rethinkingembeddingcouplingpretrained), XLM-R (conneau2019unsupervised), mBERT (devlin2019bert), and mDeBERTa (he2021debertav3) for monolingual and crosslingual experiments.
  2. 2.
[…]

### v3

@@ 3.3 Annotation Process and Guidelines
  • If the text is classified as polarized, tag its manifestations or rhetorical tactics (stereotyping/deindividuation, vilification, dehumanization, extreme language, lack of empathy, invalidation).
@@
3.4 Annotators’ Reliability
To evaluate annotation quality, we report Fleiss’ Kappa as the inter-annotator agreement (IAA) metric. As shown in Table˜1, the IAA scores vary between languages, with the majority showing moderate agreement and a few, such as "khm" and "tel" achieving good agreement. Although guidelines were standardized, their interpretation was influenced by cultural and political context, especially in languages with lower agreement, where some terms may not have direct equivalents across cultures. Latent content or sarcasm often required annotators to draw on their own socio-political knowledge, highlighting the perspectivist nature of polarization (Cabitza et al., 2023). Thus, low agreement can indicate socio-pragmatic complexity rather than error, signaling that polarization markers may not have universal meanings and that divergences can reveal inherent ambiguity in stimuli or interpretation (Aroyo and Welty, 2015). Examples illustrating such ambiguities are provided below.
Unanimous polarized: “However I got tanned in Bodrum and I don’t wait in line at the hospital.” This sentence was rated as polarized by all three annotators with high confidence. Here, the sentence refers to perceived privileges granted to the immigrant community.
Non-unanimous polarized: “Do you agree with the idea that new citizens in our country shouldn’t be allowed to vote for ten years?” This sentence did not receive consensus among annotators, leaning toward being polarized. The topic mentioned here is more complex and sensitive compared to the previous example.
3.5 Dataset Statistics
A comprehensive analysis of the dataset was conducted using the annotated labels to support systematic examination. Table 1 provides a quantitative breakdown of positive labels for each annotation task. For further details on data sources, language-wise composition, and polarization statistics, including distributions of polarized instances, types, and manifestations, see Appendix B (Tables 67 and Figure 4). Sample data is shown in Table LABEL:tab:sample_data, 10, and 11.
# Lang. PolarDetect PolarType PolarManifest
mBERT XLM-R RemBERT LaBSE twitter- roberta-hate afro-xlmr- large mBERT XLM-R RemBERT LaBSE twitter- roberta-hate afro-xlmr- large mBERT XLM-R RemBERT LaBSE twitter- roberta-hate afro-xlmr- large
1 eng 74.83 76.34 77.51 77.27 79.59 75.54 31.29 27.33 43.17 34.68 36.99 19.09 41.83 44.50 47.90 46.58 49.04 47.82
deu 65.69 69.03 69.81 68.57 65.97 59.14 49.34 48.89 54.23 53.26 49.39 51.34 43.74 45.24 47.46 47.13 44.28 45.18
urd 71.60 68.74 75.62 74.20 63.45 73.51 68.76 72.16 72.96 73.91 71.05 67.16 72.13 72.83 74.75 77.09 72.49 43.00
ben 79.03 83.68 82.74 82.50 36.81 83.04 24.61 16.59 10.14 26.82 10.14 13.62 21.08 24.12 5.05 25.83 9.59 5.64
hin 76.37 75.06 45.98 78.23 61.80 65.49 61.40 63.58 73.30 69.97 38.83 57.87 53.71 69.12 72.69 70.13 63.58 62.07
ori 41.72 41.72 71.14 77.60 42.42 68.73 19.44 28.14 32.92 42.27 19.23 29.47 8.37 4.84 2.78 24.00 5.77 9.47
nep 85.04 84.04 87.48 88.26 69.86 76.09 45.75 48.78 70.53 66.42 32.69 49.16 55.15 53.92 58.11 60.19 39.19 35.55
pan 69.98 66.50 73.57 74.91 61.03 62.28 29.59 28.22 40.88 38.73 27.32 29.08 42.07 40.07 45.80 48.35 38.57 43.31
spa 70.36 71.27 75.24 74.99 71.96 74.24 58.22 54.54 65.01 59.34 57.78 58.64 41.17 43.34 50.63 48.04 43.76 46.45
ita 54.35 34.50 60.93 56.40 53.00 47.78 21.48 21.83 25.99 26.13 21.08 23.47
rus 70.20 71.61 77.39 74.77 42.33 74.80 31.97 21.93 47.33 40.94 15.81 36.44
pol 74.45 74.54 77.96 77.26 68.45 67.62 31.63 31.29 46.84 41.55 24.65 24.48
fas 80.09 78.34 77.50 83.50 54.64 80.89 46.69 47.21 54.18 52.46 27.72 27.75 35.53 34.42 41.35 39.51 24.72 34.78
2 hau 82.97 47.19 75.91 82.09 78.32 81.82 21.44 15.75 15.53 21.58 16.68 2.76 19.97 16.43 21.44 20.55 18.29 22.16
arb 78.24 78.43 81.12 81.23 68.37 82.20 49.49 50.43 58.89 55.89 33.40 56.16 49.94 53.05 55.53 56.76 43.21 56.30
amh 42.45 68.83 72.60 76.43 49.46 58.28 29.32 39.48 24.27 47.09 28.19 28.43 41.21 26.94 42.47 51.16 41.28 45.75
3 zho 83.96 86.45 87.08 86.44 72.49 85.83 57.39 47.83 69.35 63.12 34.05 45.17 40.50 37.91 39.09 46.10 27.72 38.87
mya 84.10 82.44 84.33 86.07 70.35 82.31 42.62 34.54 25.55 55.06 26.05 42.52
4 khm 47.58 47.58 76.41 73.70 47.58 68.38 23.06 50.12 65.79 58.60 15.89 17.20 14.54 29.77 35.60 34.28 14.57 14.26
5 swa 76.54 76.02 78.17 79.04 75.93 78.77 36.39 33.20 43.35 40.18 35.63 38.46 55.69 54.32 56.21 56.51 53.79 54.82
6 tel 84.86 79.63 88.93 88.93 65.16 78.62 41.16 32.36 43.68 42.56 30.77 36.73 37.87 21.55 39.79 39.24 29.65 35.83
7 tur 69.80 74.47 70.94 75.02 65.42 73.06 36.55 43.92 51.03 48.41 30.76 31.33 36.92 44.83 44.48 44.92 36.83 44.36
Table 2: Average macro-F1(%) scores for PolarDetect, PolarType, and PolarDetect  across languages and multilingual encoders. The best and second performance scores are highlighted in blue and orange respectively.
@@ 4.1 Experimental Setup
To evaluate POLAR, we conducted baseline experiments on three polarization detection tasks: (1) classifying texts as polarized or not, (2) identifying polarization types, and (3) detecting polarization manifestations. For data splitting, we used 70% for training, 10% for validation, and 20% for testing, as summarized in Table˜7. All experiments were conducted using the EncouRAGe framework proposed by  Strich et al. (2025), which provides a standardized evaluation protocol for language model benchmarking. We benchmarked SLMs and LLMs. The list of evaluated models are listed below and Appendix C:
  • Fine-tuning SLMs: We fine-tuned six SLMs, including four general-purpose models: mBERT (Devlin et al., 2019), XLM-R Conneau et al. (2020), RemBERT (Chung et al., 2021), and LaBSE (Feng et al., 2022). In addition, we evaluated two models build on social media and multilingual training corpus: twitter-roberta-hate (Antypas and Camacho-Collados, 2023), a RoBERTa-based encoder specialized for hate-speech detection, and AfroXLMR-large-76L (Adelani et al., 2023), an XLM-R variant optimised for African and other low-resource languages.
  • Evaluating LLMs in Zero- and Few-shot Settings: We evaluated large language models in zero- and few-shot settings, including Qwen2.5-7B-Instruct, Qwen-3-8B, LLaMA-3.1-8B-Instruct, Ministral-3-14B-Instruct-2512, Gemma-3-27B-IT, GPT-4.1-Nano, and GPT-OSS-120B. For brevity, we refer to them as Qwen2.5, Qwen3, LLaMA3.1, Mistral3, Gemma3, GPT4.1 and GPT-OSS respectively. The exact models used are stated in the Appendix section C. The prompts used for LLM zero-shot and few-shot settings are shown in Appendix F.
# Lang. PolarDetect PolarType PolarManifest
Qwen2.5 Qwen3 LLaMA3.1 Mistral3 Gemma3 GPT4.1 GPT-OSS Qwen2.5 Qwen3 LLaMA3.1 Mistral3 Gemma3 GPT4.1 GPT-OSS Qwen2.5 Qwen3 LLaMA3.1 Mistral3 Gemma3 GPT4.1 GPT-OSS
1 eng 71.19 69.42 69.17 70.67 79.84 77.19 77.43 72.45 66.76 69.08 73.79 74.37 69.59 77.73 63.33 62.38 60.62 63.91 64.18 63.39 68.33
deu 60.02 58.76 61.53 61.34 73.86 70.89 69.72 66.61 61.26 64.17 67.31 71.71 63.63 70.04 57.74 60.10 56.82 58.71 57.93 59.41 65.18
urd 40.97 48.46 58.55 51.26 75.88 72.02 70.42 42.03 44.24 49.27 45.11 50.65 46.92 49.32 42.20 50.32 52.14 54.74 53.22 48.44 52.46
ben 65.62 59.75 68.90 69.83 80.85 76.15 80.28 68.96 66.10 65.71 71.05 74.04 72.64 77.13 58.47 54.54 51.23 50.47 51.89 56.99 60.34
hin 43.94 36.53 48.52 45.67 67.13 59.04 58.96 65.32 54.48 59.96 68.80 75.12 69.73 71.68 48.09 53.62 55.74 60.50 58.43 53.04 58.53
ori 53.35 50.14 49.30 42.38 65.03 58.39 67.20 65.87 64.74 65.11 48.61 75.81 66.68 71.29 55.64 55.22 54.27 49.76 59.46 57.88 61.66
nep 49.03 53.48 65.01 61.84 85.28 74.16 82.68 69.31 67.29 65.09 73.78 80.95 72.30 82.29 62.67 67.08 61.88 70.01 73.34 69.59 76.30
pan 40.69 40.42 49.80 41.64 71.95 71.67 65.92 56.51 54.20 56.98 54.21 68.20 61.73 66.19 51.10 54.90 53.20 55.22 56.73 58.48 61.94
spa 66.64 68.25 70.42 70.04 71.43 73.64 76.91 67.69 66.07 61.82 69.05 69.81 65.93 72.18 60.00 59.86 57.46 57.30 59.84 60.34 66.36
ita 66.24 63.20 67.52 67.36 73.72 72.08 74.60 68.48 64.68 63.59 70.19 73.27 68.47 74.70
rus 73.77 71.01 71.12 72.88 70.57 72.85 74.36 70.26 66.33 57.78 73.43 69.88 68.47 77.94
pol 57.15 58.33 63.46 57.25 68.88 65.42 76.94 70.56 62.91 65.26 68.87 75.88 70.42 79.46
fas 32.17 28.44 40.10 29.71 54.79 48.71 51.61 59.53 52.96 63.23 54.93 71.17 60.74 66.33 58.62 55.68 56.52 55.40 56.63 58.98 62.93
2 hau 51.40 49.25 55.37 50.44 46.36 52.18 55.02 55.41 53.65 50.69 54.59 54.04 56.84 59.03 51.01 47.28 43.09 43.56 43.66 48.43 48.33
arb 65.22 58.82 70.02 65.47 80.70 77.72 78.87 63.22 62.77 62.62 66.17 71.22 64.39 72.25 62.74 64.57 62.64 67.04 67.97 65.02 73.93
amh 25.16 26.97 39.14 23.50 74.47 50.58 57.00 52.85 51.55 61.67 47.26 77.01 61.20 67.31 50.78 53.99 56.51 53.43 59.77 55.98 61.52
3 zho 58.20 59.70 70.57 62.62 82.48 78.51 80.52 77.38 73.32 68.54 75.32 80.63 74.76 81.15 69.99 67.72 58.16 64.35 65.82 70.44 77.90
mya 40.63 48.31 51.68 37.72 78.40 58.41 59.96 52.65 55.76 58.48 49.23 67.78 58.09 59.48
4 khm 10.56 8.94 11.04 8.65 13.32 11.25 12.53 50.51 46.30 57.32 45.32 55.21 48.71 49.43 47.96 50.04 50.14 49.28 49.30 48.66 48.50
5 swa 40.50 46.76 57.42 52.29 65.64 62.00 66.15 61.31 57.52 61.63 62.87 69.49 66.34 68.81 52.94 54.04 54.63 55.11 56.89 56.22 60.41
6 tel 42.99 44.23 43.02 36.82 42.97 45.23 43.51 52.81 53.88 55.82 52.24 57.88 52.70 50.95 50.78 54.96 54.08 55.68 55.29 56.59 52.13
7 tur 64.11 58.86 66.68 66.23 76.69 74.54 78.97 67.43 63.41 62.19 70.38 75.41 70.27 77.49 60.38 61.54 60.84 62.94 65.08 62.73 71.20
Table 3: F1-Macro resulting from the zero-shot LLM experiments with the POLAR dataset. The highest value per language is highlighted in blue.